# On DR-semigroups satisfying the ample conditions

### v1

@@ 1 Preliminaries
1.1 Partial operations and categories
First, some terminology for partial operations. Let $S$ be a set with partial binary operation $\circ$. We here say that $(S,\circ)$ is a partial semigroup if $x\circ(y\circ z)$ is defined if and only if $(x\circ y)\circ z$ is, and then they are equal; equivalently (as is well-known), it is a semigroup with zero, made into a partial algebra by removing 00 and retaining all nonzero products. In a partial semigroup, we may unambiguously write $a\circ b\circ c$ (if it exists). A number of other definitions of partial semigroups exist in the literature, but the one used here is perhaps the most commonly used one.
Here we define a small category as follows: it is a set $C=(C,\circ,D,R)$ equipped with a partial binary operation $\circ$ and two unary operations of “domain” $D$ and “range” $R$ such that
  • $D(x)\circ x=x$, $x\circ R(x)=x$
  • $R(D(x))=D(x)$, $D(R(x))=R(x)$
  • $x\circ y$ exists if and only if $R(x)=D(y)$
@@ 1.1 Partial operations and categories
It follows easily that
  • if $x\circ y$ exists then $D(x\circ y)=D(x),R(x\circ y)=R(y)$.
This “object-free” formulation of the concept of a category (or some equivalent of it) is frequently used in algebra; see for example [2] and [4].
@@ 1.2 DR-semigroups
The term “DR-semigroup” was introduced in [9], but the concept was first defined in [4], in the form of so-called reduced $E$-semiabundant semigroups. These are semigroups with distinguished subset of idempotents $E$ such that the induced generalized Green’s relations ~$\tilde{\mathcal{R}}_{E}$ and ~$\tilde{\mathcal{L}}_{E}$ are such that each ~$\tilde{\mathcal{R}}_{E}$-class and each ~$\tilde{\mathcal{L}}_{E}$-class contains a member of $E$; it is further assumed that for all $e,f\in E$, $e=ef\;\Leftrightarrow\;e=fe$, and this is sufficient (but not necessary) to ensure that the member of $E$ in each ~$\tilde{\mathcal{R}}_{E}$-class and each ~$\tilde{\mathcal{L}}_{E}$-class is unique. This gives rise to two unary operations as follows: for all $s\in S$, $D(s)$ is the unique $e\in E$ for which $(s,e)\in\tilde{\mathcal{R}}_{E}$, and similarly for $R(s)$ in terms of ~$\tilde{\mathcal{L}}_{E}$. One can then equationally characterise the resulting biunary semigroups as those satisfying the following laws:
@@
Let $S$ be a DR-semigroup. If $ef=fe$ for all $e,f\in D(S)$, then we say that $D(S)$ commutes, and $D(S)$ is easily seen to be a semilattice in this case. (If $D(S)$ commutes and $e,f\in D(S)$, then $D(ef)=D(ef)D(e)=D(ef)e=D(fe)e=D(fe)fe=fe=ef$, so $ef=fe\in D(S)$.)
@@ 2 Introducing the ample conditions for DR-semigroups
If $x\leq_{r}y$ for some $x,y\in S$, then $x=D(x)y$ and $D(x)\leq D(y)$, so $x=yR(D(x)y)$, and $R(D(x)y)\leq R(y)$ by Law (DR5) for DR-semigroups, so $x\leq_{l}y$. The converse follows dually.
Next, for $e,f\in D(S)$, by (2) in Lemma 1.1 and the ample conditions, $D(ef)=D(ef)D(e)=D(ef)e=eD(f)=ef$, and so $ef=efe$. Dually, using $R$ we obtain $fe=efe$, and so $ef=fe$.
@@
Proof.    Suppose $D(S)\subseteq B(S)$. Then for all $e,f\in D(S)$, $ef=D(ef)e=D(ef)$, and dually $ef=R(ef)$, so $ef=D(ef)=D(ef)e=efe=eR(fe)=R(fe)=fe$. Hence $D(S)$ commutes.
Conversely, suppose $D(S)$ commutes. Then for all $e,f\in D(S)$, $D(ef)=ef$ and so $ef=efe=D(ef)e$, and dually, $ef=fR(ef)$, so $e\in B(S)$. $\Box$
@@ 3.1 Example: strongly order-preserving partial functions on a quasiordered set
Suppose $(X,\leq)$ is a quasiordered set. We say a partial function $f$ on $X$ is strongly order-preserving if, for all $x,y\in\mbox{dom}(f)$, $x\leq y\;\Leftrightarrow\;f(x)\leq f(y)$. Define $P_{SO}(X)$ to be the subset of all strongly order-preserving partial functions on $X$.
@@
Define domain and range operations on $P_{SO}(X)$ as follows: for $f\in P_{SO}(X)$, $D(f)$ is the restriction of the identity map to the down-set generated by $\mbox{dom}(f)$, that is, the set $\{y\in X\mid y\leq x\mbox{ for some }x\in\mbox{dom}(f)\}$. Define $R(f)$ dually using $\mbox{im}(f)$.
It follows from Lemma 12.1 in [3] that $(P_{SO}(X),\cdot,D,R)$ is a DR-semigroup in which $D(P_{SO}(X))$ commutes, which moreover is translucent in the sense of that reference, meaning that $se=D(se)s$ for all $s\in P_{SO}(X)$ and $e\in D(P_{SO}(X))$; this is nothing but the left ample condition. In fact we have the following.
@@
Such examples need not satisfy the congruence conditions. Suppose $X=\{1,2\}$, ordered as usual, let $f=\{(1,1)\}$ and let $g=\{(2,2)\}$, so that in $P_{SO}(X)$, $D(g)=I_{X}$, the identity function on $X$. Then $D(fg)=D(\emptyset)=\emptyset$, whereas $D(fD(g))=D(f)=f\neq D(fg)$.
3.2 Example: binary relations under double demonic composition
Consider the set $Rel(X)$ of binary relations on the non-empty set $X$. For $s\in Rel(X)$ and $a\in X$, we define $as=\{b\in X\mid(a,b)\in s\}$, and we define $sa$ dually.
It is well-known that $(Rel(X),\cdot,D,R)$ is an Ehresmann semigroup if $\cdot$ is relational composition, with $D$ given by
$D(s)=\{(x,x)\mid x\in X,x\in\mbox{dom}(s)\}$
and dually for $R(s)$. A related semigroup structure replaces usual “angelic” composition on $Rel(X)$ by so-called demonic composition $\star$ given by
$(x,y)\in s\star t\;\Leftrightarrow\;(x,y)\in st\>\&\>xs\subseteq\mbox{dom}(t).$
Thus $s\star t$ is a “domain restriction” of $st$, where $st$ denotes the usual, or “angelic”, composition of $s,t$.
Demonic composition is also a generalization of composition of partial functions. Like composition, demonic composition is associative and $(Rel(X),\star,D,R)$ is a DR-semigroup in which $D(S)$ commutes, and indeed is a left restriction semigroup with range in the sense of [10], meaning that it is a DR-semigroup in which $(Rel(X),\star,D)$ a left restriction semigroup. The right congruence condition $R(st)=R(R(s)t)$ is not satisfied; however, $S$ satisfies the cat-semigroup and left ample conditions.
We can obtain a DR-semigroup satisfying the ample conditions from $Rel(X)$ by applying the domain constraint used in the definition of demonic composition on the range side as well. Thus for $s,t\in Rel(X)$, we define $s\diamond t$ as follows:
$(x,y)\in s\diamond t\;\Leftrightarrow\;(x,y)\in st\>\&\>xs\subseteq\mbox{dom}(t)\>\&\>ty\subseteq\mbox{im}(s).$
We call this new operation double demonic composition of binary relations. Unlike demonic composition, this operation is left/right symmetric. It generalizes composition of injective partial functions but not composition of partial functions in general.
Lemma 3.2
Let $X$ be a non-empty set. For $s,t\in Rel(X)$, $(x,y)\in s\diamond t$ if and only if $(x,y)\in st$, $x\in\mbox{dom}(s\diamond t)$ and $y\in\mbox{im}(s\diamond t)$.
Proof.    Suppose $x,y\in X$ are such that $(x,y)\in st$, $x\in\mbox{dom}(s\diamond t)$ and $y\in\mbox{im}(s\diamond t)$. Since $x\in\mbox{dom}(s\diamond t)$, it follows that $xs\subset\mbox{dom}(t)$; dually, $ty\subseteq\mbox{im}(s)$. Hence, $(x,y)\in s\diamond t$. The converse is immediate. $\Box$
We then obtain the following.
Theorem 3.3
For a non-empty set $X$, $(Rel(X),\diamond,D,R)$ is a DR-semigroup satisfying the ample and cat-semigroup conditions but neither congruence condition in general.
Proof.    We begin with associativity of $\diamond$. Suppose $s,t,u\in Rel(X)$, and that $(a,b)\in s\diamond(t\diamond u)$ for some $a,b\in X$. Then $(a,b)\in s(t\diamond u)\subseteq stu$, so $(a,c)\in s,(c,d)\in t,(d,b)\in u$ for some $c,d\in X$, and moreover $as\subseteq\mbox{dom}(t\diamond u)$ and $(t\diamond u)b\subseteq\mbox{im}(s)$.
Now $as\subseteq\mbox{dom}(t\diamond u)$, so $as\subseteq\mbox{dom}(t)$. Also, $(t\diamond u)b\subseteq\mbox{im}(s)$, and $td\subseteq tub=(t\diamond u)b$ (since $d\in ub\subseteq\mbox{im}(t)$), so $td\subseteq\mbox{im}(s)$. We already saw that $as\subseteq\mbox{dom}(t)$ and $(a,d)\in st$, so we may deduce that $(a,d)\in s\diamond t$, so $(a,b)\in(s\diamond t)u$. Now $a(s\diamond t)=ast$, but $as\subseteq\mbox{dom}(t\diamond u)$, so $ast\subseteq\mbox{dom}(u)$, and $ub\subseteq\mbox{im}(t)$. But $tub=(t\diamond u)b\subseteq\mbox{im}(s)$, so by definition, $ub\subseteq\mbox{im}(s\diamond t)$. Hence, $(a,b)\in(s\diamond t)\diamond u$.
This shows that $s\diamond(t\diamond u)\subseteq(s\diamond t)\diamond u$. The reverse inclusion follows dually, establishing associativity of $\diamond$.
Note that if $\mbox{im}(s)=\mbox{dom}(t)$, or if $s,t$ are injective partial functions, then $s\diamond t=st$. Hence for all $s\in Rel(X)$, $D(s)\diamond s=s=s\diamond R(s)$. Of course, $D(R(s))=R(s)$, $R(D(s))=D(s)$. Also, certainly $\mbox{dom}(D(s\diamond t))\subseteq\mbox{dom}(D(s))$, and because $D(s\diamond t),D(s)$ are injective partial functions,
$D(s\diamond t)\diamond D(s)=D(s\diamond t)D(s)=D(s\diamond t)=D(s)D(s\diamond t)=D(s)\diamond D(s\diamond t)$
for all $s,t\in Rel(X)$. Hence, $(Rel(X),\diamond,D,R)$ is a DR-semigroup.
Now if $s,t\in Rel(X)$ and $R(s)=D(t)$, then $\mbox{im}(s)=\mbox{dom}(t)$, so $D(s\diamond t)=D(st)=D(sD(t))=D(sR(s))=D(s)$, and dually, $R(s\diamond t)=R(t)$. So $(Rel(X),\diamond,D,R)$ satisfies the cat-semigroup condition.
Finally, we show the ample conditions. Pick $t\in Rel(X)$ and $e\in D(Rel(X))$. Then $e\diamond t=(e\diamond t)\diamond R(e\diamond t)=e\diamond(t\diamond R(e\diamond t))\subseteq t\diamond R(e\diamond t)$. Conversely, suppose $(a,b)\in t\diamond R(e\diamond t)$. Then $(a,b)\in t$ and $at\subseteq\mbox{im}(e\diamond t)$. Now $b\in at$, so $b\in\mbox{im}(e\diamond t)$, so $b\in\mbox{im}(t)$ and for all $c\in X$ for which $(c,b)\in t$, $c\in\mbox{dom}(e)$. So, because $(a,b)\in t$, we have $a\in\mbox{dom}(e)$. Hence, $(a,b)\in et$. Since $a\in\mbox{dom}(e)$ and $a\in\mbox{dom}(t)$, we also have that $a\in\mbox{dom}(e\diamond t)$. Since $(a,b)\in et$, $a\in\mbox{dom}(e\diamond t)$ and $b\in\mbox{im}(e\diamond t)$, by Lemma 3.2, we have $(a,b)\in e\diamond t$. So $t\diamond R(e\diamond t)\subseteq e\diamond t$, completing the proof of the left ample condition. The right ample condition follows dually.
Let $x,y\in X$ with $x\neq y$, and consider $s=\{(x,y),(y,y)\}\in Rel(X)$. Then $D(s\diamond s)=D(\emptyset)=\emptyset$, but $D(s)=\{(x,x),(y,y)\}$ and so $D(s\diamond D(s))=D(s)\neq\emptyset$. Hence, the left congruence condition is not satisfied in general. Dually for the right congruence condition. $\Box$
Note that the operation of converse on binary relations, , makes $(Rel(X),\diamond,\>^{\prime}\>)$ into a semigroup with involution. When restricted to the subsemigroup of injective partial functions, both $(Rel(X),\diamond,D,R,\>^{\prime}\>)$ and $(Rel(X),\cdot,D,R,\>^{\prime}\>)$ yield the symmetric inverse semigroup with $s^{\prime}$ the inverse of $s$, $D(s)=ss^{\prime}=s\diamond s^{\prime}$, and dually for $R(s)$.

### v2

@@ 1 Preliminaries
1.1 Partial algebras
First, some terminology for partial operations. Let $S$ be a set with partial binary operation $\circ$. We here say that $(S,\circ)$ is a partial semigroup if $x\circ(y\circ z)$ is defined if and only if $(x\circ y)\circ z$ is, and then they are equal; equivalently (as is well-known), it is a semigroup with zero, made into a partial algebra by removing 0 and retaining all nonzero products. In a partial semigroup, we may unambiguously write $a\circ b\circ c$ (if it exists). A number of other definitions of partial semigroups exist in the literature, but the one used here is perhaps the most commonly used one.
Here we define a (small) partial category $(C,\circ,D,R)$ as follows: it is a set $C$ equipped with a partial binary operation $\circ$ and two unary operations of “domain” $D$ and “range” $R$ such that
  1. (PC1)
    for all $x\in C$, $D(x)\circ x$ and $x\circ R(x)$ both exist and equal $x$
  2. (PC2)
    for all $x,y\in C$, if $x\circ y$ exists then $R(x)=D(y)$
  3. (PC3)
@@ 1.1 Partial algebras
It is a category if it satisfies the following strengthening of Law (PC2) above:
  • for all $\in C, xyx\circ y$ exists if and only if $R(x)=D(y)$.
It follows easily that in the partial category $C$,
  • for all $x\in C$, $R(D(x))=D(x)$, $D(R(x))=R(x)$, and
  • for all $x,y\in C$, if $x\circ y$ exists then $D(x\circ y)=D(x),R(x\circ y)=R(y)$.
This is because, for all $x\in C$, $D(x)\circ x$ exists and so $R(D(x))=D(x)$, and so if $x\circ y$ exists then so does $D(x\circ y)\circ(x\circ y)$ whence so must $D(x\circ y)\circ x$, and so $D(x\circ y)=R(D(x\circ y))=D(x)$; dually, $D(R(x))=R(x)$ and $R(x\circ y)=R(y)$.
Such an “object-free” formulation of the concept of a category is frequently used in algebra; see for example [2] and [5]. The more general partial category notion is due to Stein, who was led to define the notion of a category with partial composition in [8]. Those are the “with objects” form of the partial categories just defined.
We shall see that partial categories arise naturally from DR-ample semigroups, but can also be used to construct them. Examples are also easily obtained as subsets of small categories.
Example 1
[…]
Proposition 1.1
If $C$ is a small category and $P\subseteq C$ is saturated, then $(P,\cdot,D,R)$ is a partial category.
Proof.   First note that if $s\in P$ then because $D(s)\circ s=s$, also $D(s)\in P$, and similarly, $R(s)\in P$. So $D(s)\cdot s=s=s\cdot R(s)$. Of course, if $x,y\in P$ and $x\cdot y$ exists in $P$, then so does $x\circ y$ in $C$, and so $R(x)=D(y)$. Finally, if $x,y,z\in P$ and $(x\cdot y)\cdot z$ exists in $P$, then $x\circ(y\circ z)=(x\circ y)\circ z$ exists in $C$ and lies in $P$, so $y\circ z\in P$, and so $y\cdot z$ exists in $P$, and similarly for $x\cdot(y\cdot z)$ which equals $x\circ(y\circ z)=(x\circ y)\circ z=(x\cdot y)\cdot z$. $\Box$
A rich family of examples arising in this way arises by letting $C$ be the small posetal category $C$ obtained from the preordered set $(X,\leq)$. In the current object-free formulation, we can take $C=\{(x,y)\in X\times X\mid x\leq y\}$, with $(x,y)\circ(u,v)=(x,v)$ if and only if $y=u$ and otherwise it is undefined, with $D((x,y))=(x,x)$ and $R((x,y))=(y,y)$. Then it is easily checked that $P\subseteq C$ is saturated if and only if, whenever $(x,z)\in P$ with $x\leq y\leq z$, it must be that $(x,y),(y,z)\in P$. For example, let $X$ be the totally ordered set of integers and define $C=\{(x,y)\in X\times X\mid x\leq y\}$ as above. If we let $P_{n}=\{(x,y)\in C\mid y-x<n\}$ for some positive integer $n$, $P_{n}$ is evidently saturated, and so defining $(x,y)\cdot(y,z)=(x,z)$ but only if zx<nz-x<n and otherwise undefined makes $(P_{n},\cdot,D,R)$ into a partial category which is evidently not a category. For example, $(2,4),(4,6)\in P_{3}$ and $(2,4)\circ(4,6)=(2,6)$ in $C$ but the result is not in $P_{3}$, so $(2,4)\cdot(4,6)$ does not exist, even though $R((2,4))=(4,4)=D((4,6))$.
Another kind of example comes from within the path category (free category) $C$ of a directed graph, if we let $C_{n}$ consist of all paths of length at most $n>0$; clearly $C_{n}$ is saturated within $C$, but $C_{n}$ is not a subcategory of $C$.
@@ 1.2 DR-semigroups
The term “DR-semigroup” was introduced in [10], but the concept was first defined in [5], in the form of so-called reduced $E$-semiabundant semigroups. These are semigroups with distinguished subset of idempotents $E$ such that the induced generalized Green’s relations $\tilde{\mathcal{R}}_{E}$ and $\tilde{\mathcal{L}}_{E}$ are such that each $\tilde{\mathcal{R}}_{E}$-class and each $\tilde{\mathcal{L}}_{E}$-class contains a member of $E$; it is further assumed that for all $f\in E, e=e$$ef\;\Leftrightarrow\;e=fe$, and this is sufficient (but not necessary) to ensure that the member of $E$ in each $\tilde{\mathcal{R}}_{E}$-class and each $\tilde{\mathcal{L}}_{E}$-class is unique. This gives rise to two unary operations as follows: for all $s\in S$, $D(s)$ is the unique $e\in E$ for which $(s,e)\in\tilde{\mathcal{R}}_{E}$, and similarly for $R(s)$ in terms of $\tilde{\mathcal{L}}_{E}$. One can then equationally characterise the resulting biunary semigroups as those satisfying the following laws:
@@
Let $S$ be a DR-semigroup. If $ef=fe$ for all $e,f\in D(S)$, then we say that $D(S)$ commutes, and $D(S)$ is easily seen to be a semilattice in this case. (If $D(S)$ commutes and $e,f\in D(S)$, then D(ef)=D(ef)D(e)=D(ef)e=D(fe)e=D(fe)fe=fe=efD(ef)=D(ef)D(e)=D(ef)e=D(fe)e=D(fe)fe=fe=ef, so $ef=fe\in D(S)$.)
@@
Let $C$ be a partial category such that $D(C)$ is partially ordered, and let $P(C)$ consist of all subsets of $C$. (Evidently, $C$ could be a category, or the partial order on $D(C)$ could be equality.) Because $C$ is a partial semigroup, $P(C)$ can obviously be made into a semigroup by defining, for all $A,B\in P(C)$,
$AB=\{a\circ b\mid a\in A,b\in B\}.$
It is shown in Example 1 in [6] that if $C$ is a small category and the partial order on $D(C)$ is taken to be equality, then $(P(C),\cdot,D,R)$ is an Ehresmann semigroup, generalizing the fact that the set of binary relations on a set gives an Ehresmann semigroup under relational composition, domain and range. More generally, we have the following.
Proposition 1.4
Proof.   Certainly $(P(C),\cdot)$ is a semigroup.
Of course, the partial order on $D(C)$ can be taken to be equality. In that case, $D,R$ are defined exactly as for the category case given in Example 3 in [6].
@@ 2 Introducing the ample conditions for DR-semigroups
If $x\leq_{r}y$ for some $x,y\in S$, then $x=D(x)y$ and $D(x)\leq D(y)$, so $x=yR(D(x)y)$, and $R(D(x)y)\leq R(y)$ by (2) in Lemma 1.2, so $x\leq_{l}y$. The converse follows dually.
Next, for $e,f\in D(S)$, by (DR4) and the ample conditions, $D(ef)=D(ef)D(e)=D(ef)e=eD(f)=ef$, and so $ef=efe$. Dually, using $R$ we obtain $fe=efe$, and so $ef=fe$.
@@
Proof.   Suppose $D(S)\subseteq B(S)$. The argument that $D(S)$ commutes is very similar to the proof given in the third part of Proposition 2.1.
Conversely, suppose $D(S)$ commutes. Then $S$ is a left C-semigroup in the sense of [3] and so $D(S)$ forms a semilattice by Proposition 1.2 there, so for all $e,f\in D(S)$, $D(ef)=ef$ and so $ef=efe=D(ef)e$, and dually, $ef=fR(ef)$, so $e\in B(S)$. $\Box$
@@ 3 Examples of DR-semigroups satisfying the ample conditions
3.1 Example: closure operators on sets
Let $X$ be a non-empty set equipped with a closure operator $C$ on its subsets. When equipped with intersection $\cap$, the power set $2^{X}$ of $X$ is a semigroup, which is a DR-semigroup in which $D(T)=R(T)=C(T)$ for all $T\subseteq X$. Moreover, it trivially satisfies the ample conditions, since $(2^{X},\cap)$ is a semilattice, but generally does not satisfy the congruence conditions.
@@ 3.2 Example: strongly order-preserving partial functions on a quasiordered set
Throughout this subsection, suppose $(X,\leq)$ is a quasiordered set.
For any $T\subseteq X$, we define its closure $C(T)$ to be the down-set generated by $T$, that is, the set $\{y\in X\mid y\leq x\mbox{ for some }x\in T\}$. It is well-known that $C$ is an additive closure operator.
We say a partial function $f$ on $X$ is strongly order-preserving if, for all $x,y\in\mbox{dom}(f)$, $x\leq y\;\Leftrightarrow\;f(x)\leq f(y)$. Define $P_{SO}(X)$ to be the subset of all strongly order-preserving partial functions on $X$.
@@
Define domain and range operations on $P_{SO}(X)$ as follows: for $f\in P_{SO}(X)$, $D(f)$ is the restriction of the identity map to $C(\mbox{dom}(f))$, the down-set generated by $\mbox{dom}(f)$. Define $R(f)$ dually using $\mbox{im}(f)$. It follows from Lemma 12.1 in [4] that $(P_{SO}(X),\cdot,D,R)$ is a DR-semigroup in which $D(P_{SO}(X))$ commutes, which moreover is translucent in the sense of that reference, meaning that $se=D(se)s$ for all $s\in P_{SO}(X)$ and $e\in D(P_{SO}(X))$; this is nothing but the left ample condition. In fact we have the following.
@@
Such examples need not satisfy the congruence conditions since they contain a copy of the power set example given earlier: the power set of $X$ arises as the restrictions of the identity to subsets of $X$ and then $D,R$ agree with $C$ as defined above on this copy of the power set of $X$.
3.3 Examples obtained from partial categories, including injective partial functions
Proposition 3.2
Proof.   Suppose $A\in B(C)$. Pick $a,b\in A$. Suppose $D(a)\leq D(b)$. Now let $E=\{e\in D(C)\mid e\leq R(b)\}$. Then $E=D(E)$, and so because $A\in B(C)$, $AE=D(AE)A$. Hence, as $b=b\circ R(b)\in AE$, we have $b\in D(AE)A$, so $b=e\circ c$ for some $e\in D(AE),c\in A$, meaning that $e=D(c)$, so $b=D(c)\circ c=c$, and so $e=D(c)=D(b)$. Since $D(a)\leq D(b)$, we have $D(a)\in D(AE)$ and therefore $a=D(a)\circ a\in D(AE)A=AE$. Hence, $a=a^{\prime}\circ f$ for some $a^{\prime}\in A$ and $f\in E$, and so $a=a^{\prime},f=R(a)\in E$ and so by definition of $E$, $R(a)\leq R(b)$. Dualising gives the converse.
Proposition 3.3
Again, the partial order on $D(C)$ can be taken to be equality. The greater generality used here is to enable a representation theorem for DR-semigroups satisfying the ample conditions in Section 5.
An example that is not a restriction semigroup may be obtained by letting $C$ consist of the posetal category obtained from the integers described after Example 1, so that $C=\{(x,y)\mid x,y\in{\mathbb{Z}},x\leq y\}$, with $D((x,y))=(x,x),R((x,y))=(y,y)$ and $(x,y)\circ(u,v)=(x,v)$ exists if and only if $y=u$. Then as noted there, defining $P_{n}=\{(x,y)\mid y-x<n\}$ for some positive integer $n$, setting $(x,y)\cdot(y,z)$ to be undefined if $z-x\geq n$ and otherwise to equal $(x,z)$ makes $S=(S,\cdot,D,R)$ into a partial category. Equip $D(P_{n})=D(C)$ with the partial order of equality. Then the elements of ${\mathcal{P}I}(S)$ may be identified with the injective partial functions $f$ defined on $\mathbb{Z}$} for which $0\leq xf-x<n$ for all $x\in\mbox{dom}(f)$, with $fg$ considered undefined at $x$ if $xfg-x\geq n$. For example, for $f=\{(1,3),(3,7)\},g=\{(3,5),(7,9)\}\in S_{5}$, we have $fg=\{(1,5)\}$, so $D(fg)=\{(1,1)\}$, whereas $D(fD(g))=D(fR(f))=D(f)=\{(1,1),(3,3)\}$, witnessing the failure of the congruence conditions.
Other examples can be obtained from saturated subsets of categories as in Example 1.